OpenStudy (anonymous):

Can someone show me how to solve this please? (I don't know how to get rid of the fraction) |3/4+2x|=2

OpenStudy (anonymous):

|3/4+2x|=2 3/4+2x=2 or 3/4+2x=-2 -3/4 -3/4 -3/4 -3/4 2x=2-3/4 or 2x=-2-3/4 2x=2-.75 or 2x=-2-.75 2x=1.25 or 2x-2.75 -- --- -- --- 2 2 2 2 x=.75 or x=1.375

OpenStudy (anonymous):

Is there another possible way without the decimal?

OpenStudy (anonymous):

|3/4+2x|=2 3/4+2x=2 or 3/4+2x=-2 -3/4 -3/4 -3/4 -3/4 2x=2-3/4 or 2x=-2-3/4 2x=1 1/4 or 2x = -2 3/4 -- ----- -- ------ 2 2 2 2 x=3/4 or -11/8

OpenStudy (anonymous):

Thank you very much c: